OpenStudy (anonymous):

determine wheather each order pair is a soultion for 2y-x=10 (2,-6) and (0,5)

OpenStudy (anonymous):

plug in and check (2,-6) 2(-6)-2 = ? (0,5) 2*5 - 0 =?

OpenStudy (anonymous):

Yeah and if they both equal 10 then they are solutions, and if either one of them do not then they are not solutions to that equation.

OpenStudy (anonymous):

how do yow show work for that qustion

OpenStudy (anonymous):

2y-x=10

OpenStudy (anonymous):

by plugging in

OpenStudy (anonymous):

You just plug in your first order pair to the x and y and do the same for the second order pair. That is showing your work.

OpenStudy (anonymous):

(2,-6) 2y -x =10 2(-6)-2 = ? if you get 10, it is a solution if not, its not a solution (0,5) 2y -x =10 2*5 - 0 = ? if you get 10, it is a solution if not, its not a solution

OpenStudy (anonymous):

@race_mom_23 did you get it?

OpenStudy (anonymous):

so the anwser ar -14 and 10

OpenStudy (anonymous):

so which one of these 2 satisfy the equation 2y - x = 10

OpenStudy (anonymous):

(0,5)

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

i have some more qustions?

OpenStudy (anonymous):

show that the orders pair(0,-3) and (2,1) are solution of the equation 2x-y=3. Then use the graph of equation to determine another solution.

OpenStudy (anonymous):

you just do the same thing as before. And then just pick a point that lies on the line for another point.

OpenStudy (anonymous):

so i take 3+-3

OpenStudy (anonymous):

Well to determine a point you need to solve the equation for y and then graph it and then just pick a point that is on that line.

OpenStudy (anonymous):

im still lost sorry